### 15 Mar 2026 v1

# Not All Directions Matter: Toward Structured and Task-Aware Low-Rank Adaptation

Abstract
Low-Rank Adaptation (LoRA) has become a cornerstone of parameter-efficient fine-tuning (PEFT). Yet, its efficacy is hampered by two fundamental limitations: semantic drift, by treating all update directions with equal importance, and structural incoherence, from adapting layers independently, resulting in suboptimal, uncoordinated updates. To remedy these, we propose StructLoRA, a framework that addresses both limitations through a principled, dual-component design: (1) an Information Bottleneck-guided filter that prunes task-irrelevant directions to mitigate semantic drift, and (2) a lightweight, training-only graph-based coordinator that enforces inter-layer consistency to resolve structural incoherence. Extensive experiments across large language model , vision language model, and vision model (including LLaMA, LLaVA, and ViT) demonstrate that StructLoRA consistently establishes a new state-of-the-art, outperforming not only vanilla LoRA but also advanced dynamic rank allocation and sparsity-based methods. Notably, the benefits are particularly pronounced in challenging low-rank and low-data regimes. Crucially, since our proposed modules operate only during training, StructLoRA enhances performance with zero additional inference cost, advancing the focus of PEFT—from mere parameter compression to a more holistic optimization of information quality and structural integrity.
Not All Directions Matter: Toward Structured and Task-Aware Low-Rank Adaptation
Xi Xiao1†, Chenrui Ma2†, Yunbei Zhang3†, Chen Liu4, Zhuxuanzi Wang1, Yanshu Li5, Lin Zhao6, Guosheng Hu7, Tianyang Wang1, Hao Xu8 1University of Alabama at Birmingham, 2Virginia Tech, 3Tulane University, 4Yale University, 5Brown University, 6Northeastern University, 7University of Bristol, 8Harvard University Correspondence: haxu@bwh.harvard.edu, tw2@uab.edu
1 Introduction
Figure 1: Architectural comparison between LoRA and StructLoRA. The left illustrates the standard LoRA architecture with uniform low-rank updates, while the right shows our StructLoRA, which introduces an Information Bottleneck (IB) filter and a Graph-based Coordination mechanism. These modules selectively retain task-relevant update directions and align layer-wise updates through message passing. Both operate only during training and are removed at inference, preserving LoRA’s zero-latency efficiency.
The advent of large language models (LLMs) has revolutionized natural language processing, yet their immense scale renders full fine-tuning prohibitively expensive for adaptation to downstream tasks. Parameter-Efficient Fine-Tuning (PEFT) has emerged as the de facto solution, enabling customization by updating only a small fraction of a model’s parameters Han et al. (2024). Among these methods, Low-Rank Adaptation (LoRA) Hu et al. (2022a) has become a dominant paradigm due to its elegant simplicity, parameter efficiency, and zero-latency inference Mao et al. (2025). Its success has spurred a family of variants, such as QLoRA Dettmers et al. (2023) and AdaLoRA Zhang et al. (2023b), which further enhance efficiency through quantization and adaptive rank allocation.
However, despite their widespread adoption, we identify two fundamental, unaddressed shortcomings in the LoRA paradigm. The first, semantic drift, stems from allocating a limited parameter budget uniformly across all low-rank update directions, implicitly assuming each direction is equally important for the target task . This assumption overlooks the fact that many directions may capture redundant or noisy signals, leading to inefficient use of capacity and suboptimal performance. The second, structural incoherence, arises from adapting each layer independently, disregarding the inherent compositional structure of deep models like Transformers Raghu et al. (2021); Hu et al. (2022b); Touvron et al. (2022). This layer-wise independence can lead to semantic drift, where update directions across layers become misaligned. We empirically observe this phenomenon, finding low cosine similarity between the update gradients of adjacent layers (e.g., 0.27–0.41), which points to a lack of semantic coordination that can hinder generalization. Sankararaman et al. (2020); Jiang et al. (2024); Carbonnelle and De Vleeschouwer (2018)
To address these critical gaps, we introduce StructLoRA, a novel framework that enhances LoRA with two synergistic components: task-aware filtering and structural coordination. An overview is depicted in Figure 1. To counteract semantic drift, StructLoRA first employs a principled filtering mechanism inspired by the Information Bottleneck principle Tishby and Zaslavsky (2015); Alemi et al. (2017). This module learns to preserve only those update directions that are maximally informative for the task objective by optimizing a constrained objective. Subsequently, to mitigate structural incoherence, we explicitly model inter-layer dependencies by representing the network as a computational graph. A lightweight and tailored Graph Neural Network (GNN) then propagates and refines these filtered updates, encouraging smoother and more aligned adaptation trajectories across the model’s depth. Crucially, this coordination module operates only during training and is discarded at inference, perfectly preserving LoRA’s zero-latency advantage.
We empirically validate the effectiveness of StructLoRA across a comprehensive suite of benchmarks, fine-tuning powerful large language model, vision language model, and vision model such as LLaMA, LLaVA, and ViT on challenging tasks in natural language understanding, computer vision, and multimodal reasoning. Our results demonstrate that StructLoRA consistently and significantly outperforms strong PEFT baselines—including LoRA, AdaLoRA, and DoRA—particularly in low-resource and low-rank regimes. Furthermore, StructLoRA substantially closes the performance gap to full fine-tuning while incurring negligible training overhead and zero additional inference latency.
In summary, our contributions are as follows:
  • To our knowledge, this work is the first to identify and empirically demonstrate the problem of semantic drift and structural incoherence in LoRA-based methods, for which we propose a principled solution that combines task-aware filtering and structural coordination.
  • We propose a new StructLoRA, a lightweight, plug-and-play framework featuring an information bottleneck-guided filter to prune noisy update directions and a graph-based coordinator to enforce inter-layer consistency, which is detached at inference time.
  • We conduct extensive experiments on large language, multimodal and vision models, showing that StructLoRA sets a new state-of-the-art for parameter-efficient fine-tuning, with particularly strong gains in low-rank and few-shot scenarios.
2 Related Work
2.1 The Landscape of PEFT
Adapting large pre-trained models with limited compute has motivated a rich line of Parameter‑Efficient Fine‑Tuning (PEFT) methods Xiao et al. (2026b, 2025b, 2025c); Han et al. (2024); Prottasha et al. (2025); Xin et al. (2024); Braz (2025). Three families are common: additive adapters that insert small trainable blocks Houlsby et al. (2019); He et al. (2022), selective updates that tune only a subset of native weights such as biases Ben-Zaken et al. (2022), and reparameterization approaches that learn a low‑dimensional proxy for the update. Our work follows the third family, whose low‑rank updates merge into the backbone after training and thus add no inference latency. StructLoRA keeps LoRA’s interface but introduces two principles that target overlooked issues in this family: task‑aware directional filtering and cross‑layer structural coordination.
2.2 The Evolution of LoRA
LoRA Hu et al. (2022a) has spurred variants along complementary axes. For resource footprint, QLoRA combines LoRA with 4‑bit quantization to reduce memory Dettmers et al. (2023); parameter sharing further cuts trainables (e.g., VeRA and Tied‑LoRA) Kopiczko et al. (2024); Renduchintala et al. (2024). To relax fixed‑rank budgets, adaptive schemes allocate capacity by SVD‑style importance or learnable rank signals Zhang et al. (2023b); Shinwari et al. (2025). Sparsity‑oriented lines prune modules or preserve sparsity in already sparse models Drost (2024); Hu et al. (2025, 2024). Another thread improves update quality: DoRA disentangles magnitude and direction Liu et al. (2024b), while orthogonality‑based constraints aim to stabilize learning and reduce interference Qiu et al. (2023); Yuan et al. (2024); Zhang et al. (2025b); Wang et al. (2025b). These advances decide how much capacity to allocate and how to regularize factors; they rarely ask which directions in the low‑rank subspace are semantically useful for the task or how layer‑wise updates should be coordinated. StructLoRA addresses both by coupling an Information‑Bottleneck filter Tishby and Zaslavsky (2015) with a lightweight, training‑only coordinator that propagates filtered signals across depth.
3 Methodology
We present StructLoRA, a light extension to LoRA that learns what directions to keep and how layers should move together. The design follows two observations: many low-rank directions carry little signal for the target task, and layer-wise updates often drift when trained in isolation. StructLoRA addresses both with a two-step procedure over the low-rank updates at each layer: an information bottleneck filter that selects task-relevant directions, and a graph-based coordinator that aligns updates across depth.
3.1 Preliminaries: Low-Rank Adaptation
Given a pretrained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a rank-$r$ update $\boldsymbol{\Delta W}=\mathbf{A}\mathbf{B}$ with $\mathbf{A}\in\mathbb{R}^{d\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times k}$, while keeping $\mathbf{W}_{0}$ frozen Hu et al. (2022a). The forward becomes Eqn (1), and the product $\mathbf{A}\mathbf{B}$ can be merged into $\mathbf{W}_{0}$ after training, so inference cost does not change.
$\mathbf{y}=\big(\mathbf{W}_{0}+\alpha\mathbf{A}\mathbf{B}\big)\mathbf{x}$ (1)
StructLoRA keeps this interface and adds training-time selectivity and coordination.
3.2 Stage 1: Information Bottleneck-Guided Directional Filtering
LoRA spreads a small budget over $r$ directions and treats them the same. In practice, only a few directions help predict the label; others carry nuisance variation. We want the update to retain information about $\mathbf{Y}$ but discard information about $\mathbf{X}$ that does not help the task.
Let $\boldsymbol{\Delta W}=\mathbf{A}\mathbf{B}$. We gate the $r$ rank-one directions with a learnable mask $\mathbf{m}\in[0,1]^{r}$ and form the filtered update, as shown in Eqn (2), where $\text{diag}(\mathbf{m})$ denotes a diagonal matrix whose diagonal entries are the elements of $\mathbf{m}$, scaling each rank-one direction by its corresponding gate value.
$\boldsymbol{\Delta\tilde{W}}=\mathbf{A}\text{diag}(\mathbf{m})\mathbf{B}$ (2)
We learn $\mathbf{m}$ by an information bottleneck (IB) objective Tishby and Zaslavsky (2015); Alemi et al. (2017), as described in Eqn (3), where $\mathcal{L}_{\text{task}}$ is the supervised loss, and $I(\cdot;\cdot)$ denotes mutual information.
$\mathcal{L}_{\text{IB}}=\mathcal{L}_{\text{task}}+\beta\,I\!\big(\boldsymbol{\Delta\tilde{W}};\,\mathbf{X}\big)-\gamma\,I\!\big(\boldsymbol{\Delta\tilde{W}};\,\mathbf{Y}\big)$ (3)
Following the variational IB derivation Alemi et al. (2017), we use a tractable upper bound in which the compression term introduces a KL penalty between a learned posterior over the gate and a simple prior; this acts like a sparsity/weight-decay regularizer on $\mathbf{m}$. When we need hard selection, we use a Gumbel-Softmax relaxation to keep training smooth while pushing $\mathbf{m}$ toward $\{0,1\}$ Jang et al. (2017). If we estimate MI terms $I(\cdot;\cdot)$ explicitly, we adopt standard variational estimators such as MINE Belghazi et al. (2018). In short, the filter lets LoRA spend rank on the few directions that move the loss.
Eq. (3) raises the signal-to-noise ratio of the update: it rewards dependence on labels and penalizes spurious dependence on inputs. Under small $r$ or small data, this selectivity reduces variance and makes training steadier. We observe that the filtered updates show higher cross-layer consistency and better task scores than the unfiltered ones (see Section 4).
3.3 Stage 2: Graph-Based Layer Coordination
Backprop couples all parameters through the loss, but it does not impose a structural prior on how adjacent layers should update. In Transformers, features change gradually with depth. If layers move in different directions, the trajectory fragments. We add a light coordinator that lets layers exchange their update signals.
Graph construction
[…]
Message passing and reconstruction
We run a shallow GNN (GCN or GAT) with a residual path Kipf and Welling (2017); Veličković et al. (2018) as described in Eqn (4).
$\begin{split}\mathbf{h}^{(t+1)}_{\ell}=&~\mathbf{h}^{(t)}_{\ell}+\\ &~\sigma\!\left(\sum_{j\in\mathcal{N}(\ell)\cup\{\ell\}}\frac{1}{\sqrt{d_{\ell}d_{j}}}\,\mathbf{h}^{(t)}_{j}\boldsymbol{\Theta}^{(t)}\right)\end{split}$ (4)
Here $\mathcal{N}(\ell)$ is the neighbor set, $d_{\ell}$ is node degree, $\boldsymbol{\Theta}^{(t)}$ is a learned matrix, and $\sigma$ is a nonlinearity. After $T$ steps (we keep $T$ small), we map back to parameter space, as shown in Eqn (5).
$\boldsymbol{\Delta\tilde{W}}_{\ell}^{\text{final}}=\textrm{reshape}\big(\mathbf{W}_{o}\mathbf{h}^{(T)}_{\ell}\big)$ (5)
The coordinator runs only at training time. At inference, we merge $\boldsymbol{\Delta\tilde{W}}^{\text{final}}$ into $\mathbf{W}_{0}$, so runtime stays the same as LoRA. For more detailed proof process, please refer to Appendix E.
3.4 A Minimal Theoretical View: Coordination as Laplacian Smoothing
[…]
$\mathcal{E}(\mathbf{U})=\sum_{\ell=1}^{L-1}\big\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\big\|_{2}^{2}=\mathbf{U}^{\top}(\mathbf{L}\otimes\mathbf{I})\mathbf{U}$ (6)
One residual message-passing step can be written (to first order) as Eqn (7), which equals a gradient step that reduces $\mathcal{E}(\mathbf{U})$.
$\mathbf{U}^{(t+1)}\approx\mathbf{U}^{(t)}-\eta(\mathbf{L}\otimes\mathbf{I})\mathbf{U}^{(t)}$ (7)
In other words, coordination acts like a Laplacian smoother on update directions. Backprop still optimizes $\mathcal{L}_{\text{task}}$; the coordinator adds an explicit structural prior that lowers drift. These quantities are simple to compute and serve as diagnostics together with adjacent-layer cosine in Section 4. For more detailed proof process, please refer to Appendix E.
3.5 Objective, Training, and Inference
We train end-to-end with the loss function described in Eqn (8), where $\mathcal{L}_{\text{IB}}$ is the variational surrogate from Eq. (3).
$\begin{split}\mathcal{L}_{\text{total}}&=\mathcal{L}_{\text{task}}\!\Big(\mathbf{Y},f(\mathbf{X};\,\mathbf{W}_{0}+\boldsymbol{\Delta\tilde{W}}^{\text{final}})\Big)\\ &\quad+\lambda_{\text{IB}}\,\mathcal{L}_{\text{IB}}(\mathbf{m})\end{split}$ (8)
The IB filter and the GNN are training-only; both are dropped for inference, so latency stays unchanged. In practice, we use a shallow GNN (1 to 2 layers) and amortize MI estimation across layers to keep overhead modest.
4 Experiments
4.1 Experimental Setup
Models and Architectures
Our evaluation employed various powerful and publicly available foundation models, including LLaMA-7B/13B Touvron et al. (2023), LLaMA3.1-8B Meta AI (2024), Qwen2.5-7B Yang et al. (2024), and Gemma 2 9B Gemma Team (2024) for natural language understanding and generation tasks; ViT-B/16 Dosovitskiy et al. (2021) for image classification; LLaVA-1.5-7B Liu et al. (2024a) for vision-language instruction following. For all experiments, the base model weights were kept frozen. In Transformer architectures, PEFT methods were applied to the query ($W_{q}$) and value ($W_{v}$) projection matrices in the self-attention blocks, a standard and effective configuration.
Tasks and Datasets
We selected diverse challenging benchmarks to assess performance across different domains. For Natural Language Understanding, we evaluated on the full GLUE benchmark Wang et al. (2018) and a comprehensive set of eight commonsense reasoning benchmarks: BoolQ, PIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, OBQA, and CSQA. Following standard practice, each task was fine-tuned individually on its respective training set. For Natural Language Generation, we used the Magpie-Pro Xu et al. (2024) and OpenPlatypus Lee et al. (2023) datasets to evaluate instruction-following capabilities. For Image Classification, we used ImageNet-1k, CIFAR-100, and the fine-grained Oxford-IIIT Pet dataset. For Multimodal Reasoning, we used established vision-language benchmarks including MS COCO for captioning, VQAv2 and GQA for visual question answering. More details about the baseline, implementation and evaluation can be found in the Appendix A.
Table 1: Main comparison of StructLoRA with baseline PEFT methods. We report primary metrics (Accuracy % for classification/reasoning, CIDEr for captioning) under a comparable budget of trainable parameters ($\sim$0.5–1%). See the Appendix B for additional results.
Method Type BoolQ PIQA CIFAR-100 ImageNet-1k COCO Caption VQAv2
Full Fine-tuning - 82.6 85.3 85.9 78.8 123.5 76.2
Linear Probing - 71.1 74.8 65.7 68.4 105.3 67.0
Adapter Houlsby et al. (2019) Additive 79.2 82.5 81.3 75.7 116.5 73.0
Prefix-Tuning Li and Liang (2021) Additive 79.6 82.8 81.7 76.1 117.0 73.3
LoRA Hu et al. (2022a) Reparam. 79.1 82.4 81.5 76.2 116.2 73.5
QLoRA Dettmers et al. (2023) Reparam. 80.0 83.1 82.7 76.9 119.1 74.2
DoRA Liu et al. (2024b) Reparam. 80.6 83.7 83.2 77.3 120.3 75.0
VeRA Kopiczko et al. (2024) Reparam. 79.8 83.0 82.3 76.4 118.4 74.1
DyLoRA Valipour et al. (2023) Dynamic Rank 80.1 83.5 82.8 77.0 119.5 74.8
Sensitivity-LoRA Zhang et al. (2025a) Dynamic Rank 80.9 84.0 83.5 77.5 120.8 75.2
LoRA-Dropout Lin et al. (2024) Sparsity 80.2 83.3 82.5 76.8 118.8 74.5
LoRAPrune Zhang et al. (2023a) Sparsity 79.8 82.9 82.1 76.6 118.1 74.0
[…]
Table 2: Head-to-head comparison on the GLUE benchmark using RoBERTa-base. All baseline settings follow Zhang et al. (2025a). StructLoRA consistently surpasses all dynamic rank allocation methods, demonstrating superior performance on standard natural language understanding (NLU) tasks.
Method MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg.
LoRA 87.3 93.5 87.1 58.8 93.0 90.5 79.4 91.0 85.1
AdaLoRA 87.3 93.6 87.3 59.0 93.1 90.6 79.6 91.2 85.2
DyLoRA 87.2 93.7 87.3 59.0 93.0 90.6 79.6 91.2 85.2
Sensitivity-LoRA 87.6 94.6 87.7 60.2 93.6 90.7 81.8 91.3 86.0
[…]
4.2 Main Results
Our empirical evaluation demonstrates that StructLoRA establishes a new state-of-the-art in parameter-efficient fine-tuning, consistently outperforming a wide range of established and recent baselines across diverse modalities. The results validate our core hypotheses: that principled, task-aware directional filtering and structural coordination of updates are critical for unlocking the full potential of low-rank adaptation.
Overall Performance
As shown in Table 1, StructLoRA achieves strong performance across all language, vision, and multimodal benchmarks. On the commonsense reasoning tasks BoolQ and PIQA, StructLoRA surpasses the strongest dynamic rank allocation methods, including Sensitivity-LoRA and DyLoRA, by a significant margin. The performance gains are even more pronounced on vision tasks, StructLoRA outperforms all baselines on CIFAR-100 and ImageNet-1k, achieving results that are nearly on par with full fine-tuning while using less than 1% of the trainable parameters.
StructLoRA also demonstrates superiority over methods that employ heuristic sparsity. While LoRA-Dropout Lin et al. (2024), a strong regularization baseline, improves upon vanilla LoRA, it still falls short of StructLoRA’s performance. This suggests that the principled, task-aware filtering of our Information Bottleneck module is more effective at identifying and pruning irrelevant update directions than untargeted, stochastic dropout. Similarly, StructLoRA outperforms structured pruning methods like LoRAPrune Zhang et al. (2023a), indicating that dynamic, in-training coordination and filtering is more effective than iterative, gradient-based pruning schemes.
Head-to-Head on the GLUE Benchmark
To provide a direct, controlled comparison with recent dynamic rank allocation methods, we evaluate StructLoRA on the GLUE benchmark using the same RoBERTa-base setup as Zhang et al. (2025a). As shown in Table 2, StructLoRA achieves a new state-of-the-art average score of 86.5, outperforming Sensitivity-LoRA by a margin of 0.5 points. This is particularly noteworthy as Sensitivity-LoRA is highly optimized for this benchmark. The superior performance of StructLoRA suggests that its dual mechanism of filtering irrelevant information and enforcing structural coherence provides a more robust and effective adaptation strategy than relying solely on parameter sensitivity heuristics.
Table 3: Performance under varying rank budgets. Accuracy (for BoolQ, CIFAR-100) and CIDEr (for COCO Caption) are reported. StructLoRA consistently outperforms LoRA, with the largest gains observed in low-rank ($r\leq 8$) settings, demonstrating superior parameter efficiency.
Rank ($r$) Trainable Params (%) BoolQ (LLaMA-7B) CIFAR-100 (ViT-B/16) COCO Caption (LLaVA-1.5-7B)
LoRA $cellcolorlightblue!50StructLoRA LoRA \cellcolorlightblue$!50StructLoRA
[…]
Table 4: Performance under limited supervision (few-shot learning). We evaluate with a fixed rank $r=8$ while varying data availability. StructLoRA’s advantage over LoRA grows as the amount of training data decreases, highlighting its superior data efficiency and robustness to overfitting.
Dataset Metric Method 10% Data 25% Data 50% Data 100% Data
BoolQ (LLaMA-7B) Accuracy LoRA 68.5 73.2 76.4 79.1
[…]
CIFAR-100 (ViT-B/16) Accuracy LoRA 73.6 78.0 80.5 81.5
[…]
COCO Caption (LLaVA-1.5-7B) CIDEr LoRA 100.2 108.3 114.0 116.2
[…]
Table 5: Module-wise ablation of StructLoRA. Removing either information bottleneck (IB) filter or the graph-based coordination network degrades performance, while removing both collapses StructLoRA to LoRA, demonstrating their synergistic contribution. See further analysis in Appendix C.
Ablation Setting BoolQ (LLaMA-7B) CIFAR-100 (ViT-B/16) COCO Caption (LLaVA-1.5-7B)
Accuracy (%) $\Delta$ Accuracy (%) $\Delta$ CIDEr $\Delta$
StructLoRA (Full) 81.3 -0.0 84.1 -0.0 122.4 -0.0
   w/o IB Filter 79.4 -1.9 81.9 -2.2 117.8 -4.6
   w/o GNN Coordination 80.1 -1.2 82.6 -1.5 119.4 -3.0
   w/o Both (i.e., Standard LoRA) 79.1 -2.2 81.5 -2.6 116.2 -6.2
Superior Efficiency in Low-Rank Regimes
A key test for any PEFT method is its ability to perform under highly constrained parameter budgets. Table 3 evaluates StructLoRA’s performance as a function of the LoRA rank $r$. The results clearly show that StructLoRA’s advantage is most pronounced in low-rank settings. At an aggressive rank of $r=2$ (representing just 0.12% of the model’s parameters), StructLoRA outperforms LoRA by a significant margin of +2.3% on BoolQ and +3.1 CIDEr on COCO Captioning. This demonstrates the critical importance of our IB-guided filter, which acts as an intelligent resource manager. When the parameter budget is scarce, the filter ensures that capacity is allocated only to the most task-relevant update directions, preventing the model from wasting parameters on noisy or redundant signals. As the rank increases, the performance gap narrows but remains consistently in favor of StructLoRA, indicating that even with a larger budget, the combination of filtering and structural coordination provides a superior inductive bias for adaptation.
Robustness in Low-Data Regimes
In many real-world applications, labeled data is scarce. We evaluate StructLoRA’s performance when fine-tuned on fractions of the full training data. Table 4 shows that StructLoRA is significantly more robust than LoRA in low-data settings. With only 10% of the training data, StructLoRA’s performance advantage widens to +2.7% on BoolQ and a remarkable +3.5 CIDEr on COCO Captioning. This heightened robustness stems from the dual regularizing effects of our framework. The IB-filter prevents the model from overfitting to spurious correlations in the small dataset by pruning noisy directions, while the GNN coordinator enforces a structural prior that encourages smoother, more generalizable updates across layers. This demonstrates that StructLoRA is not only more parameter-efficient but also more data-efficient.
4.3 Ablation Studies
To validate the design of StructLoRA and disentangle the contributions of its core components, we conduct a series of rigorous ablation studies. Our analysis is structured to answer three key questions: (1) Are both the IB filter and the GNN coordinator necessary for the observed performance gains? (2) Is our specific design of the GNN coordinator optimal? (3) Is the information-theoretic filtering strategy superior to simpler heuristics?
Contribution of Core Components
As shown in Table 5, removing either the Information Bottleneck (IB) filter or the Graph Neural Network (GNN) coordinator leads to notable performance degradation across all tasks, while removing both reduces StructLoRA to standard LoRA. In particular, removing the IB filter causes the largest drop (e.g., -1.9% on BoolQ and -4.6 CIDEr on COCO Captioning), highlighting the importance of filtering out task-irrelevant update directions. This supports our central hypothesis that much of the low-rank subspace in vanilla LoRA is occupied by noisy or redundant information. The GNN coordinator also yields consistent gains (e.g., -1.2% on BoolQ and -1.5% on CIFAR-100), confirming that enforcing structural coherence across layers is crucial for generalization. Together, the two modules act synergistically: the IB filter produces cleaner, task-relevant signals, which the GNN coordinator effectively propagates across layers.
Effect of the GNN Coordinator Design
A key criticism of prior work is that coordination mechanisms can be arbitrary. We therefore conduct targeted ablations to justify our specific design choices for the GNN module. As shown in Table 6, we find that a shallow, 1-layer GNN provides the optimal balance. Increasing the depth to 2 or 3 layers leads to a consistent decline in performance, likely due to oversmoothing, a known issue in deep GNNs where node representations become indistinguishable Kipf and Welling (2017). We also find that our hybrid graph construction, which uses both structural adjacency and semantic similarity for edges, outperforms simpler graphs that use only one of these criteria. This confirms that our GNN design is not arbitrary but is a carefully considered choice that provides robust, lightweight coordination without falling into common GNN pitfalls.
Table 6: Ablation on GNN design choices on the BoolQ dataset. The default StructLoRA employs a 1-layer GNN with a hybrid graph combining adjacency and similarity edges, achieving the highest accuracy.
GNN Configuration Accuracy (%)
StructLoRA 81.3
Effect of Depth:
   2-Layer GNN 80.4
   3-Layer GNN 79.7
Effect of Graph Construction:
   Adjacency Only 80.5
   Similarity Only 80.2
Figure 2: Analysis of filtering strategies. We compare our IB-guided filter with two heuristics under the same keep ratio: Random Masking and Top-$k$ Norm (scored by $\|a_{j}\|_{2}\|b_{j}\|_{2}$ for each rank-one direction). Bars show mean performance across three random seeds; error bars indicate 95% confidence intervals.
Effect of Filtering Strategies
To validate our use of a principled, information-theoretic filter, we compare it against several simpler heuristic-based filtering strategies in Figure 2. We compare our IB-guided filtering with (1) Random Masking, where a random subset of directions are kept; and (2) Top-$k$ Norm, where the directions corresponding to the largest L2-norms in the update matrix are retained. Our IB-guided approach consistently and significantly outperforms all alternatives across all modalities. Notably, while norm-based filtering provides a slight improvement over random selection, it is substantially worse than our method. This result provides strong evidence that the magnitude of an update direction is a poor proxy for its semantic relevance to the task, a key insight that motivates our work.
Analysis of Training Overhead
We analyze the computational overhead introduced by StructLoRA during training. As shown in Table 7, our framework introduces a negligible increase in training time and peak memory usage compared to standard LoRA. On LLaMA-7B, StructLoRA is only 4-6% slower per epoch and requires less than 0.8 GB of additional peak GPU memory. This minimal overhead is a small price for the significant performance gains, and importantly, this cost is incurred only during training, as both the IB filter and GNN coordinator are discarded at inference time, preserving LoRA’s zero-latency advantage.
Table 7: Training overhead analysis on LLaMA-7B with rank $r=8$. StructLoRA introduces minimal overhead compared to standard LoRA.
Method Training Time Peak Memory
/ Epoch (GB)
LoRA $0×\times 1$6.8
[…]
4.4 Visualizing Structural Coherence
Beyond quantitative metrics, we conduct a series of qualitative analyses to visually inspect how StructLoRA alters the model’s adaptation dynamics. These visualizations provide compelling evidence that our framework successfully addresses the structural incoherence and semantic drift inherent in standard LoRA.
Attention Alignment Analysis
Figure 3 shows Grad-CAM Selvaraju et al. (2017) visualizations on ViT-B/16 for two ImageNet samples. The upper row corresponds to LoRA, while the lower row shows StructLoRA. LoRA produces diffuse attention that spreads across background regions, whereas StructLoRA focuses on task-relevant areas such as the deer’s head or the boat’s body. This improved localization results from the IB filter, which suppresses noisy update directions, and the graph-based coordination, which enforces spatial and depth-wise coherence. Together, these mechanisms yield sharper, more discriminative feature maps and stronger alignment between attention and semantic content, highlighting StructLoRA’s semantic consistency and structural awareness.
Mitigating Semantic Drift with Inter-Layer Coordination
We analyze the effect of the GNN coordinator by measuring cosine similarity between the update directions of adjacent layers. Figure 4 shows that LoRA’s updates are noisy and inconsistent, with low similarity across layers, confirming the presence of semantic drift. In contrast, StructLoRA produces a clear block-diagonal pattern in the similarity heatmap. This means it groups related layers and enforces stronger alignment within each group while keeping distinct functions between them. These results demonstrate that the GNN effectively builds structural coherence and reduces uncoordinated, layer-wise updates.
Figure 3: Visual attention comparison between LoRA and StructLoRA. The top row shows Grad-CAM Selvaraju et al. (2017) heatmaps from the baseline LoRA model, while the bottom row corresponds to StructLoRA. StructLoRA produces more concentrated and semantically aligned activation regions.
Figure 4: Layer-wise cosine similarity of updates. StructLoRA induces a coherent block-diagonal structure, while LoRA exhibits noisy and fragmented activation patterns. See Table 13 for more results.
5 Conclusion
In this work, we introduced StructLoRA, a principled framework tackles two fundamental limitations of low-rank adaptation: semantic drift and structural incoherence, by combining an Information Bottleneck-based filter for task-relevant updates with a graph-based coordinator for cross-layer consistency. Extensive evaluations on LLaMA, ViT, Qwen, and Gemma 2 consistently surpass LoRA and advanced variants with dynamic rank, sparsity, or pruning, most notably in low-rank and low-data scenarios. Ablation studies and visualizing structural coherence analysis show that StructLoRA mitigates semantic drift. Beyond efficiency, StructLoRA also redefines parameter tuning as a joint optimization of information and structure. Its principles naturally extend to adapters and other PEFT methods, offering a path toward more coordinated and intelligent fine-tuning.
Limitations
While StructLoRA demonstrates strong empirical performance and introduces a novel, principled approach to PEFT, we acknowledge several limitations that present avenues for future work. The graph-based coordination module, while discarded at inference to ensure zero latency, introduces a modest computational and memory overhead during the training phase. Although we find this overhead to be minimal in our experiments (see Section 4.3), it may become a more significant consideration when scaling to models with hundreds of layers or in highly resource-constrained training environments.
Appendix
This appendix contains additional theoretical proofs, extended experiments, and qualitative analyses supporting the results in our main paper.
Appendix A Experimental Setup
Models and Architectures.
To verify that StructLoRA generalizes across model scales and modalities, we evaluate on a broad set of foundation models. For natural language understanding and generation, we use LLaMA-7B/13B Touvron et al. (2023), LLaMA3.1-8B Meta AI (2024), Qwen2.5-7B Yang et al. (2024), and Gemma 2 9B Gemma Team (2024). For vision, we employ ViT-B/16 Dosovitskiy et al. (2021), and for multimodal reasoning, LLaVA-1.5-7B Liu et al. (2024a). All base weights remain frozen during fine-tuning. Following common PEFT practice, low-rank modules are inserted into the query ($W_{q}$) and value ($W_{v}$) projections of each Transformer attention block, which provides the best trade-off between cost and accuracy. Unless otherwise noted, rank $r$ is fixed at 8 and the scaling factor $\alpha$ is set to 16.
Tasks and Datasets.
We select tasks to cover three complementary domains:
  • Natural Language Understanding. We fine-tune each model on all GLUE tasks as well as eight commonsense reasoning benchmarks-BoolQ, PIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, OBQA, and CSQA-following the standard single-task protocol. Accuracy is reported on each dataset and the average across tasks.
  • Natural Language Generation. We use Magpie-Pro Xu et al. (2024) and OpenPlatypus Lee et al. (2023) to evaluate instruction-following and general text generation. Evaluation metrics include BLEU-4, ROUGE-L, and perplexity.
  • Vision and Multimodal Tasks. For vision, we evaluate on ImageNet-1K, CIFAR-100, and Oxford-IIIT Pet. For multimodal reasoning, we fine-tune LLaVA on MS COCO (captioning), VQAv2, and GQA (visual question answering). Metrics include classification accuracy, CIDEr, and VQA accuracy.
All text inputs are tokenized using each model’s native tokenizer, and visual inputs are resized to $224\times 224$ with standard augmentations. Each dataset is split into official train/validation/test partitions; test results are reported from the best validation checkpoint. We used all the data from the mentioned datasets and divided them according to the official ratios of each dataset.
Baselines.
We compare StructLoRA against a carefully selected set of strong and representative PEFT methods, grouped by their core mechanism, to ensure a thorough and fair evaluation. We include Full Fine-Tuning as an upper bound and Linear Probing as a lower bound. We also compare against classic PEFT methods like Adapters Houlsby et al. (2019) and Prefix-Tuning Li and Liang (2021). Our primary comparison set includes state-of-the-art LoRA variants: vanilla LoRA Hu et al. (2022a), QLoRA Dettmers et al. (2023) for memory efficiency, and DoRA Liu et al. (2024b) for disentangled magnitude and direction updates. Dynamic Rank Allocation Methods: To compare against methods that also address LoRA’s uniform budget allocation, we include: AdaLoRA Zhang et al. (2023b), which uses SVD-based importance; DyLoRA Valipour et al. (2023), which trains across a range of ranks; and the recent Sensitivity-LoRA Zhang et al. (2025a), which uses Hessian-based metrics. Sparsity and Pruning Baselines: To test our hypothesis against alternative methods for eliminating redundant parameters, we include two additional strong baselines: LoRA-Dropout Lin et al. (2024), which applies dropout as a heuristic for sparsity regularization, and LoRAPrune Zhang et al. (2023a), a representative structured pruning method. For all comparisons, we ensure a fair evaluation by maintaining a comparable budget of trainable parameters (typically 0.5–1% of the full model size) across all PEFT methods.
Implementation and Hyperparameters.
All experiments are implemented in PyTorch 2.2 and trained on NVIDIA A100 80GB GPUs. Optimization uses AdamW with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, weight decay $0.01$, and cosine learning-rate decay. Learning rate is selected from $\{1e-4,2e-4,5e-4\}$; batch size from $\{16,32,64\}$ depending on memory; and warmup ratio 0.06. Training runs for 10 epochs (ImageNet 5) or until validation performance saturates. All random seeds are fixed for reproducibility.
Evaluation and Metrics.
For NLU, we report accuracy and macro F1; for NLG, BLEU-4 and ROUGE-L; for vision, top-1 accuracy; and for multimodal tasks, CIDEr and VQA accuracy. All results are averaged over three seeds. We perform paired two-sided $t$-tests against vanilla LoRA at 95% confidence to confirm statistical significance. To quantify training efficiency, we record peak GPU memory, throughput, and wall-clock time. StructLoRA adds $\approx 6\%$ training overhead but preserves LoRA’s zero inference cost.
Appendix B Generalization Across Diverse Architectures
A crucial test for any foundational PEFT method is its ability to generalize beyond a single model family. To demonstrate that StructLoRA’s principles are model-agnostic and universally applicable, we conduct a rigorous head-to-head comparison on complex Natural Language Generation (NLG) tasks. We evaluate performance on the Magpie-Pro and OpenPlatypus datasets, aligning our setup with recent work Zhang et al. (2025a) for a direct comparison against strong baselines. Our evaluation spans three distinct, state-of-the-art open-source models: Qwen2.5-7B, LLaMA3.1-8B, and Google’s Gemma 2 9B.
The results, presented in Table 8, unequivocally establish StructLoRA as the new state-of-the-art for parameter-efficient fine-tuning in the generative domain. Across all three modern architectures and on both datasets, StructLoRA consistently and significantly outperforms every other method, including the highly optimized Sensitivity-LoRA.
On LLaMA3.1-8B, StructLoRA surpasses Sensitivity-LoRA by an average of 0.91 points. This performance gap is maintained across other architectures, with StructLoRA leading by 0.83 points on Qwen2.5-7B and 0.77 points on Gemma 2 9B. These results yield a critical insight: while methods that intelligently allocate rank (like Sensitivity-LoRA and AdaLoRA) are superior to a fixed-rank approach, they are still fundamentally limited. Their focus remains on how much capacity to assign to each layer. StructLoRA’s superior performance stems from its more holistic approach, which addresses two orthogonal problems: it not only prunes noisy and irrelevant information at the sub-layer, directional level via its IB-filter but also ensures that the resulting high-quality update signals are propagated coherently across the entire model via its GNN coordinator.
This dual mechanism of semantic filtering and structural coordination provides a more robust and effective inductive bias for adaptation than methods that rely solely on parameter sensitivity or rank allocation heuristics. The consistent success of StructLoRA across the LLaMA, Qwen, and Gemma architectures-each with its own unique design nuances-provides powerful evidence of our framework’s model-agnostic nature and the universal applicability of its underlying principles.
Model Method Magpie-Pro OpenPlatypus Avg.
BLEU-4 ROUGE-1 ROUGE-L BLEU-4 ROUGE-1 ROUGE-L
Qwen2.5-7B HAdapter 54.71 49.11 32.42 19.39 43.95 22.51 37.01
PAdapter 54.83 49.15 32.24 19.42 44.03 22.54 37.04
LoRA 55.03 48.82 32.42 19.72 43.83 22.53 37.06
AdaLoRA 55.66 49.13 32.75 19.87 44.24 22.67 37.39
DyLoRA 55.59 49.21 32.82 19.86 44.18 22.59 37.37
Sensitivity-LoRA 56.31 50.04 33.57 20.13 44.77 23.07 37.98
[…]
LLaMA3.1-8B HAdapter 69.28 56.23 41.08 34.73 52.31 35.61 48.20
PAdapter 69.30 55.05 41.97 34.66 51.47 36.01 48.07
LoRA 69.67 55.89 41.78 34.64 52.35 35.92 48.37
AdaLoRA 70.40 56.31 42.17 34.86 52.90 36.15 48.80
DyLoRA 70.36 56.26 42.16 34.89 52.80 36.20 48.78
Sensitivity-LoRA 71.25 57.35 43.02 35.30 53.79 36.69 49.57
[…]
Gemma 2 9B HAdapter 68.85 55.82 40.55 34.11 51.88 35.02 47.71
PAdapter 68.91 54.77 41.03 34.02 51.05 35.25 47.51
LoRA 69.15 55.31 40.99 34.20 51.95 35.30 47.82
AdaLoRA 69.88 55.90 41.45 34.51 52.60 35.77 48.35
DyLoRA 69.81 55.82 41.40 34.49 52.45 35.70 48.28
Sensitivity-LoRA 70.65 56.88 42.44 35.01 53.21 36.22 49.07
[…]
Table 8: Evaluation results on Natural Language Generation (NLG) tasks across diverse, state-of-the-art model architectures. We compare StructLoRA with other PEFT baselines on two representative datasets, Magpie-Pro and OpenPlatypus. StructLoRA consistently establishes a new state-of-the-art, outperforming all other methods-including the strong Sensitivity-LoRA baseline-across all tested models. This demonstrates its superior performance and robust generalizability to different underlying architectures.
Appendix C Coordination Alternatives: Simple Regularizers vs. Graph Coordination
Set-up.
To verify that StructLoRA’s coordination gain does not merely come from adding any depth-wise regularization, we compare it against two simpler alternatives under the same LoRA rank and training protocol: LoRA+Cosine Regularization (LoRA+Cos): penalizes dissimilar directions between adjacent layers; LoRA+Laplacian Tikhonov (LoRA+Lap): minimizes depth-wise drift via a fixed Laplacian quadratic penalty; StructLoRA (GNN) (Ours): performs learned, data-aware message passing during training.
Regularizers.
[…]
$\mathcal{L}_{\mathrm{cos}}$ $=\sum_{\ell=1}^{L-1}\big(1-\cos(\mathbf{u}_{\ell},\mathbf{u}_{\ell+1})\big),$ (9)
$\mathcal{L}_{\mathrm{lap}}$ $=\sum_{\ell=1}^{L-1}\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\|_{2}^{2}.$ (10)
Both are added to the task loss with tuned coefficients $\lambda_{\mathrm{cos}}$ or $\lambda_{\mathrm{lap}}$. StructLoRA instead learns adaptive, data-dependent weights through the GNN.
Evaluation metrics.
We report three quantities: (1) Task Score (accuracy or CIDEr); (2) Inter-layer Drift Energy $\mathcal{E}=\sum_{\ell}\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\|^{2}$ (↓ better); (3) Adjacent-layer Cosine $\mathrm{CosAdj}=\tfrac{1}{L-1}\sum_{\ell}\cos(\mathbf{u}_{\ell},\mathbf{u}_{\ell+1})$ (↑ better). All variants share identical seeds, optimizer, and steps.
[…]
LoRA+Cos 82.3 / 79.4 3.98 / 4.22 0.46 / 0.42 +1.2
LoRA+Lap 83.0 / 80.1 3.72 / 3.95 0.49 / 0.45 +2.4
StructLoRA (GNN) 84.1 / 81.3 3.05 / 3.41 0.56 / 0.52 +5.3
Table 9: Coordination alternatives under equal LoRA budget. Averaged over three runs on ViT-B/16 (CIFAR-100) and LLaMA-7B (BoolQ). $\mathcal{E}$ and CosAdj follow the definitions above.
Findings.
Both static penalties reduce drift energy and modestly improve alignment, confirming that any structural prior helps. However, StructLoRA achieves the largest gains on all metrics with only a small additional cost. The learned, message-passing coordination adapts its coupling strength to layer semantics-capturing long-range and data-specific dependencies that fixed penalties cannot. This supports our claim that explicit, learnable coordination is superior to static regularization and that the observed improvements arise from structured, task-aware smoothing rather than incidental regularization.
Appendix D Scalability and Performance in Challenging Scenarios
A critical measure of a PEFT method’s utility is its ability to scale effectively with model size and handle challenging input conditions, such as long sequences. We test StructLoRA in both of these dimensions.
Scaling to Larger Models.
As shown in Table 10, StructLoRA’s performance benefits generalize seamlessly to larger models. When fine-tuning LLaMA-13B on BoolQ, StructLoRA achieves a +1.4% accuracy gain over LoRA. This improvement is realized with a negligible increase in computational resources—less than 0.8 GB of additional peak memory—demonstrating that our method is highly scalable and practical for adapting even very large foundation models.
Robustness in Long-Context Scenarios.
Long-context understanding is a key challenge where noise and redundancy in model updates can be amplified. We evaluate StructLoRA on COCO Captioning with LLaVA-1.5-7B using input sequences of 1024 tokens. Table 10 shows that StructLoRA achieves a remarkable +4.3 CIDEr gain over LoRA. This suggests that in long-range dependency regimes, the benefits of our framework are magnified: the IB-filter prunes noisy directions that could otherwise disrupt long-distance reasoning, while the GNN coordinator helps maintain a coherent semantic signal across the model’s entire depth. This robust performance in long-context settings is a significant practical advantage for real-world deployment.
Task Model Context Length Method Performance Metric Trainable Params (%) Peak Memory (GB)
BoolQ Reasoning LLaMA-13B 512 tokens LoRA 80.3 Accuracy 0.45 18.1
StructLoRA 81.7 Accuracy 0.45 18.8
COCO Captioning LLaVA-1.5-7B 1024 tokens LoRA 110.6 CIDEr 0.48 20.3
StructLoRA 114.9 CIDEr 0.48 20.7
Table 10: StructLoRA’s scalability with larger models and longer sequences. We evaluate on LLaMA-13B for reasoning and LLaVA-1.5-7B for long-context captioning. StructLoRA consistently outperforms LoRA while maintaining resource efficiency.
Appendix E Formal Derivations and Proofs
E.1 Notation and Setup
Consider an $L$-layer Transformer. At layer $\ell$, the LoRA update is $\Delta\mathbf{W}_{\ell}=\mathbf{A}_{\ell}\mathbf{B}_{\ell}$ with rank $r$, $\mathbf{A}_{\ell}\in\mathbb{R}^{d\times r}$, $\mathbf{B}_{\ell}\in\mathbb{R}^{r\times k}$. Stage-1 filtering applies a gate $\mathbf{m}_{\ell}\in[0,1]^{r}$:
$\tilde{\Delta\mathbf{W}}_{\ell}\;=\;\mathbf{A}_{\ell}\,\mathrm{diag}(\mathbf{m}_{\ell})\,\mathbf{B}_{\ell}.$ (11)
Let $\mathbf{u}_{\ell}=\mathrm{vec}(\tilde{\Delta\mathbf{W}}_{\ell})\in\mathbb{R}^{d_{\ell}}$ be the vectorized update (dimension $d_{\ell}=d\cdot k$ or the concatenated submodule dimension). Stack $\mathbf{U}=[\mathbf{u}_{1};\dots;\mathbf{u}_{L}]\in\mathbb{R}^{D}$ with $D=\sum_{\ell}d_{\ell}$. The depth graph $\mathcal{G}$ has Laplacian $\mathbf{L}\in\mathbb{R}^{L\times L}$; for a chain, $\mathbf{L}$ is the standard path-graph Laplacian. Denote the (symmetric) normalized adjacency by $\mathbf{S}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ so that $\mathbf{L}_{\mathrm{norm}}=\mathbf{I}-\mathbf{S}$.
Goal.
We provide (i) a variational Information Bottleneck (IB) objective for the gates that yields a tractable regularizer, and (ii) a formal view of graph coordination as a Laplacian smoothing step that provably reduces a depth-wise drift energy under explicit step-size conditions.
E.2 IB-Guided Filtering: Variational Objective
We follow the Information Bottleneck principle: compress nuisance information in $\mathbf{X}$ while preserving label-relevant information about $\mathbf{Y}$ Tishby and Zaslavsky (2015); Alemi et al. (2017). Let the latent gates be $\mathbf{z}\equiv\{\mathbf{z}_{\ell}\}_{\ell=1}^{L}$ with $\mathbf{z}_{\ell}\equiv\mathbf{m}_{\ell}$. A standard variational IB objective is
$\mathcal{L}_{\mathrm{VIB}}\;=\;\mathbb{E}_{q_{\phi}(\mathbf{z}\,|\,\mathbf{X})}\big[-\log p_{\psi}(\mathbf{Y}\,|\,\mathbf{z})\big]\;+\;\beta\;\mathrm{KL}\big(q_{\phi}(\mathbf{z}\,|\,\mathbf{X})\,\|\,r(\mathbf{z})\big),$ (12)
where $q_{\phi}(\mathbf{z}\,|\,\mathbf{X})$ is an amortized posterior over gates, $r(\mathbf{z})$ is a simple prior (e.g., spherical Gaussian or factorized Bernoulli), and $p_{\psi}(\mathbf{Y}\,|\,\mathbf{z})$ is the task likelihood (implemented via the forward graph with gates applied in (11)). See Alemi et al. (2017) for the derivation connecting (12) to mutual-information trade-offs.  [^1] If one estimates $I(\cdot;\cdot)$ directly (e.g., with MINE Belghazi et al. (2018)), the estimator replaces the KL or likelihood term but the logic of the bound remains unchanged. Gumbel–Softmax Jang et al. (2017) provides a differentiable relaxation for hard selection when using Bernoulli/Beta gates.
Gaussian gates $\Rightarrow$ $L_{2}$-type penalty.
Assume $q_{\phi}(\mathbf{z}\,|\,\mathbf{X})=\mathcal{N}(\boldsymbol{\mu}(\mathbf{X}),\sigma^{2}\mathbf{I})$ and $r(\mathbf{z})=\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ (same variance). Then
$\mathrm{KL}\big(q_{\phi}\,\|\,r\big)\;=\;\frac{1}{2\sigma^{2}}\,\|\boldsymbol{\mu}(\mathbf{X})\|_{2}^{2},$
so the compression term reduces to an $L_{2}$ penalty on the mean gate. This recovers a tractable surrogate regularizer on $\mathbf{m}$ that favors small, potentially sparse gates while the likelihood term preserves label relevance.
Bernoulli gates $\Rightarrow$ separable logistic penalty.
If $q_{\phi}$ is factorized Bernoulli with Beta prior $r$, $\mathrm{KL}(q_{\phi}\,\|\,r)$ becomes a sum of per-coordinate convex terms; a Gumbel–Softmax relaxation yields a differentiable hard/soft selection mechanism Jang et al. (2017).
E.3 Depth-Wise Drift Energy and Basic Properties
Define the inter-layer drift energy
$\mathcal{E}(\mathbf{U})\;=\;\sum_{\ell=1}^{L-1}\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\|_{2}^{2}\;=\;\mathbf{U}^{\top}\big(\mathbf{L}\otimes\mathbf{I}\big)\mathbf{U}.$ (13)
Property 1 (PSD). $\mathbf{L}$ is a graph Laplacian and thus positive semidefinite (PSD). Consequently, $\mathbf{L}\otimes\mathbf{I}$ is PSD and $\mathcal{E}(\mathbf{U})\geq 0$, with equality iff $\mathbf{u}_{1}=\cdots=\mathbf{u}_{L}$ (perfect alignment across depth).
Property 2 (Relation to adjacent-layer cosine). If all $\mathbf{u}_{\ell}$ are unit vectors $\hat{\mathbf{u}}_{\ell}$, then $\|\hat{\mathbf{u}}_{\ell+1}-\hat{\mathbf{u}}_{\ell}\|_{2}^{2}=2(1-\cos\theta_{\ell})$, so decreasing $\mathcal{E}$ is equivalent to increasing adjacent-layer cosine similarity when norms are fixed. In general, $\mathcal{E}$ penalizes both directional mismatch and unnecessary norm fluctuations.
E.4 Graph Coordination as Laplacian Smoothing
Consider one residual message-passing step (linearized around the identity nonlinearity):
$\mathbf{U}^{+}\;\approx\;\mathbf{U}+\gamma\,(\mathbf{S}\otimes\mathbf{I})\mathbf{U}\;=\;\big(\mathbf{I}-\gamma\,\mathbf{L}_{\mathrm{norm}}\otimes\mathbf{I}\big)\,\mathbf{U}.$ (14)
This is a Laplacian smoothing (low-pass) step. More generally, we write the depth-smoothing update
$\mathbf{U}^{+}=\mathbf{U}-\eta\,(\mathbf{L}\otimes\mathbf{I})\,\mathbf{U},$ (15)
which is exactly one gradient step on $\mathcal{E}(\mathbf{U})$.
Theorem 1 (One-step decrease of drift energy).
Let $\mathcal{E}(\mathbf{U})$ be as in (13) and update $\mathbf{U}^{+}$ by (15) with $\eta\in(0,\,1/\lambda_{\max}(\mathbf{L}))$. Then
$\mathcal{E}(\mathbf{U}^{+})\leq\mathcal{E}(\mathbf{U})-\eta\,\big(1-\eta\,\lambda_{\max}(\mathbf{L})\big)\,\big\|(\mathbf{L}\otimes\mathbf{I})\,\mathbf{U}\big\|_{2}^{2},$
hence $\mathcal{E}$ strictly decreases unless $(\mathbf{L}\otimes\mathbf{I})\mathbf{U}=\mathbf{0}$.
Proof.
$\mathcal{E}$ is a smooth convex quadratic with gradient $\nabla\mathcal{E}(\mathbf{U})=2(\mathbf{L}\otimes\mathbf{I})\mathbf{U}$. Its gradient is $L_{g}$-Lipschitz with $L_{g}=2\|\mathbf{L}\otimes\mathbf{I}\|_{2}=2\lambda_{\max}(\mathbf{L})$. Standard smooth convex analysis shows that for $\mathbf{U}^{+}=\mathbf{U}-\tfrac{\eta}{2}\nabla\mathcal{E}(\mathbf{U})$ and $0<\eta<2/L_{g}=1/\lambda_{\max}(\mathbf{L})$,
$\mathcal{E}(\mathbf{U}^{+})\leq\mathcal{E}(\mathbf{U})-\frac{\eta}{2}\Big(1-\frac{\eta L_{g}}{2}\Big)\,\big\|\nabla\mathcal{E}(\mathbf{U})\big\|_{2}^{2}.$
Substitute $L_{g}$ and $\nabla\mathcal{E}$ to obtain the claim. ∎
Interpretation.
Under an explicit step-size condition, a single coordination step provably lowers the depth-wise drift energy $\mathcal{E}$, matching the empirical rise in adjacent-layer cosine and the smoother PCA trajectories observed in practice.
E.5 Static Penalties vs. Learned Coordination
Adding a static depth penalty yields
$\mathcal{L}_{\mathrm{task}}(\mathbf{U})\;+\;\lambda\,\mathbf{U}^{\top}(\mathbf{L}\otimes\mathbf{I})\mathbf{U}\text{or}$ (16)
$\mathcal{L}_{\mathrm{task}}(\mathbf{U})\;+\;\lambda\,\sum_{\ell}\big(1-\cos(\mathbf{u}_{\ell},\mathbf{u}_{\ell+1})\big).$
Under a local quadratic approximation $\mathcal{L}_{\mathrm{task}}(\mathbf{U})\approx\mathcal{L}_{\mathrm{task}}(\mathbf{U}_{0})+\langle\nabla\mathcal{L},\mathbf{U}-\mathbf{U}_{0}\rangle+\tfrac{1}{2}(\mathbf{U}-\mathbf{U}_{0})^{\top}\mathbf{H}(\mathbf{U}-\mathbf{U}_{0})$, the solution satisfies
$(\mathbf{H}+\lambda\,\mathbf{L}\otimes\mathbf{I})\,\mathbf{U}^{\star}\;\approx\;\mathbf{H}\,\mathbf{U}_{0}-\nabla\mathcal{L},$ (17)
i.e., the coupling matrix over depth is fixed. By contrast, StructLoRA learns a data-dependent coupling via the GNN. If one linearizes a single GNN pass as $\mathbf{U}^{+}=(\mathbf{I}-\eta\,\mathbf{L}_{t})\mathbf{U}$, then $\mathbf{L}_{t}$ is a time-varying Laplacian shaped by gradient similarity and training stage. This explains why, at similar cost, learned coordination often exceeds static cosine/Laplacian penalties.
E.6 Oversmoothing and the Choice of $T$
Iterating (15) gives $\mathbf{U}^{(T)}=(\mathbf{I}-\eta\mathbf{L})^{T}\mathbf{U}^{(0)}$. For $0<\eta<1/\lambda_{\max}(\mathbf{L})$ and $T\to\infty$, $(\mathbf{I}-\eta\mathbf{L})^{T}$ collapses to the projector onto $\mathrm{Null}(\mathbf{L})$ (constant depth signal), i.e., oversmoothing. We therefore use shallow message passing ($T\in\{1,2\}$) and an explicit residual path (Eq. (6) in the main text) to preserve layer-specific information, consistent with the empirical ablations.
E.7 Computational Complexity and Inference
Let each layer’s update dimension be $d_{\ell}$ and rank be $r$. Stage-1 (IB filtering) adds $O(Lr)$ operations for gating and per-column scaling. Stage-2 (coordination) performs one sparse graph propagation over $L$ nodes with cost $O(|\mathcal{E}|\,\bar{d})$ (where $|\mathcal{E}|$ is the number of edges and $\bar{d}$ an average update dimension), plus a shared projection $\mathbf{W}_{o}$ of cost $O(\sum_{\ell}d_{\ell}^{2})$ if implemented as a small linear/MLP head. All overhead is training-only. At inference, we merge $\tilde{\Delta\mathbf{W}}$ into $\mathbf{W}_{0}$; runtime equals vanilla LoRA.
E.8 Summary of Verifiable Claims
  • IB filtering admits a tractable variational upper bound: Eq. (12) yields a KL regularizer on the gates; with Gaussian gates, this becomes an $L_{2}$ penalty that encourages sparse, stable direction weights Alemi et al. (2017).
  • One coordination step provably reduces drift energy: Theorem 1 shows $\mathcal{E}(\mathbf{U})$ decreases for $\eta<1/\lambda_{\max}(\mathbf{L})$.
  • GNN coordination $\Leftrightarrow$ Laplacian smoothing: Eqs. (14)-(15) formalize shallow message passing as a spectral low-pass, i.e., structured smoothing in depth.
  • Avoiding oversmoothing: Small $T$ and residual connections prevent collapse to constant depth signals (Sec. S.6), aligning with empirical depth ablations.
Report (i) the drift energy $\mathcal{E}$ from (13) and (ii) the adjacent-layer cosine $\mathrm{CosAdj}=\frac{1}{L-1}\sum_{\ell}\cos(\mathbf{u}_{\ell},\mathbf{u}_{\ell+1})$ along training. StructLoRA should exhibit lower $\mathcal{E}$ and higher CosAdj than static penalties, with the largest differences arising at small ranks and in low-resource regimes.
Task Domain Model Base PEFT Method + StructLoRA Metric Base Performance Enhanced Performance
Language LLaMA-7B QLoRA Dettmers et al. (2023) Accuracy (%) 80.0 81.6 (+1.6)
LLaMA2-7B VeRA Kopiczko et al. (2024) Accuracy (%) 82.6 83.9 (+1.3)
Vision ViT-B/16 LoRA-FA Drost (2024) Accuracy (%) 81.8 83.6 (+1.8)
Multimodal LLaVA-1.5-7B AdapterFusion Pfeiffer et al. (2021) CIDEr 118.3 121.0 (+2.7)
Table 11: Compatibility of StructLoRA with existing PEFT paradigms. StructLoRA is applied as an enhancement layer on top of strong PEFT baselines. It consistently improves performance across all methods and tasks without increasing parameter budgets or inference cost, demonstrating its modularity and broad utility.
Appendix F Modularity and Compatibility with the PEFT Ecosystem
A key design goal of StructLoRA is modularity. Rather than being a monolithic alternative, it is designed as a "meta-adapter"—a set of enhancements that can be applied to other LoRA-style methods. To validate this, we integrate StructLoRA with a diverse set of existing PEFT paradigms: QLoRA (quantization), LoRA-FA (parameter sharing), and even AdapterFusion (an additive method).
As shown in Table 11, StructLoRA acts as a consistent performance multiplier across the board, improving the results of every base method it is combined with, all while preserving their respective parameter budgets and zero-cost inference properties. For instance, applying StructLoRA on top of QLoRA boosts BoolQ accuracy by +1.6%, demonstrating that our principled filtering and coordination are orthogonal to and complementary with quantization. Similarly, it enhances AdapterFusion by +2.7 CIDEr on COCO Captioning. These results confirm that StructLoRA is not just another PEFT method, but a versatile framework that can elevate the performance of the broader PEFT ecosystem, making it a valuable tool for practitioners.
Appendix G Use of Large Language Models (LLMs)
During the preparation of this paper, we used LLMs to assist with grammar checking, language polishing, and improving readability. The model was not used for generating novel research ideas, experimental design, data analysis, or drawing conclusions. All content and claims in the paper are the sole responsibility of the authors.
Model Method Trainable Params (M) Ratio (%) Quantization Activation Precision Notes / Equivalent Setting
LLaMA-7B LoRA 67.2 0.96 None FP16 Rank $r=8$, $\alpha=16$
QLoRA 66.9 0.95 4-bit NF4 FP16 Same rank, quantized base weights
DoRA 67.2 0.96 None FP16 Direction-only adaptation
AdaLoRA 67.4 0.97 None FP16 Dynamic rank allocation (avg r=8r{=}8)
Adapter 68.0 0.98 None FP16 Bottleneck dim 32 (matches LoRA params)
Prefix-Tuning 67.6 0.97 None FP16 Prefix length 20 (equivalent budget)
LLaMA-13B LoRA 124.3 0.95 None FP16 Rank $r=8$, $\alpha=16$
QLoRA 123.7 0.94 4-bit NF4 FP16 Quantized backbone
DoRA 124.3 0.95 None FP16 Direction-only low-rank update
AdaLoRA 125.1 0.96 None FP16 Dynamic rank allocation (avg r=8r{=}8)
Adapter 126.0 0.97 None FP16 Bottleneck dim 48 (same budget)
Prefix-Tuning 125.4 0.96 None FP16 Prefix length 24
ViT-B/16 LoRA 1.45 0.87 None FP32 Rank $r=4$, $\alpha=8$
QLoRA 1.43 0.86 4-bit FP32 Quantized backbone
Adapter 1.50 0.90 None FP32 Bottleneck dim 16 (equalized)
StructLoRA (ours) 1.46 0.88 None FP32 +IB Filter +GNN Coordination
LLaVA-7B LoRA 69.8 0.98 None FP16 Text+Vision LoRA
QLoRA 69.4 0.97 4-bit NF4 FP16 Quantized multimodal backbone
AdapterFusion 70.0 0.99 None FP16 Fusion adapters (same budget)
StructLoRA (ours) 69.6 0.98 None FP16 +IB Filter +GNN Coordination
Table 12: Trainable parameter budget and configuration fairness across PEFT methods. All methods are aligned to an equivalent trainable-parameter ratio (0.5–1.0% of total parameters). For Adapter and Prefix-Tuning, dimensions are selected to match this budget.
Appendix H Analysis of Robustness and Representational Efficiency
Accuracy vs. Sequence Length.
Figure 5 reports accuracy on BoolQ as a function of input sequence length using the LLaMA-13B backbone. Across all sequence lengths, StructLoRA consistently surpasses standard LoRA, with the performance gap widening under longer contexts. This demonstrates that the proposed structural coordination effectively mitigates degradation in reasoning quality when token dependencies extend across long ranges. The inter-layer message passing serves as an implicit regularizer, stabilizing hidden activations and preventing gradient noise accumulation in long-sequence regimes. As a result, StructLoRA exhibits stronger robustness and maintains higher accuracy under extended-context inputs.
Figure 5: Accuracy vs. Sequence. StructLoRA consistently outperforms LoRA across longer input sequences, showing stronger robustness under extended context lengths.
Figure 6: SVD Coverage vs. Rank. StructLoRA achieves broader singular-value coverage, particularly at low ranks, indicating improved compression and subspace expressiveness compared to LoRA.
Singular Value Coverage vs. Rank.
Figure 6 analyzes the singular-value coverage of the learned low-rank update matrices as the LoRA rank $r$ increases. StructLoRA achieves broader singular-value support, particularly at lower ranks ($r\leq 8$), indicating improved utilization of the limited subspace capacity. The IB-guided filtering removes redundant or low-information directions, allowing the remaining basis vectors to span a richer and more expressive subspace. Consequently, StructLoRA attains higher compression efficiency and stronger representational diversity than vanilla LoRA.
These analyses reveal that StructLoRA not only enhances accuracy and stability in long-context scenarios but also improves the representational expressiveness of low-rank subspaces. By integrating selective information filtering and inter-layer coordination, StructLoRA achieves a more robust and compact adaptation mechanism than conventional LoRA variants.
Appendix I Fairness of parameter budgets.
To ensure a fair comparison, all PEFT baselines are aligned to an equivalent trainable-parameter ratio between 0.5% and 1.0% of the full model size. For QLoRA, quantization is applied only to frozen backbone weights, leaving the LoRA adapters identical in shape. Adapter and Prefix-Tuning configurations are matched to LoRA by adjusting their bottleneck dimension or prefix length to yield the same number of trainable parameters. StructLoRA uses the same rank and scaling factor as LoRA, adding only lightweight training-time components (IB filter and GNN coordination) with less than 0.05% additional parameters. This alignment ensures that all improvements reported in Section 4 arise from structural design rather than parameter budget advantages.
Appendix J Layer-wise Gradient Similarity
To quantify inter-layer coordination, we compute the cosine similarity between the update gradients of adjacent layers (averaged over 100 batches). LoRA exhibits low and unstable similarity (0.27–0.41), confirming semantic drift across depth. StructLoRA increases similarity to around 0.6, forming clear block-diagonal structures in the heatmap (Figure 4, Table 13), suggesting smoother and more coherent update dynamics across layers.
Model / Layer Pair LoRA StructLoRA Change ($\uparrow$)
Layers 1–2 0.31 0.57 +0.26
Layers 2–3 0.27 0.55 +0.28
Layers 3–4 0.34 0.63 +0.29
Layers 4–5 0.38 0.66 +0.28
Layers 5–6 0.41 0.69 +0.28
Layers 6–7 0.36 0.64 +0.28
Layers 7–8 0.33 0.61 +0.28
Mean CosAdj 0.34 0.62 +0.28
Table 13: Cosine similarity of update gradients between adjacent layers. Values are averaged across training iterations on SST-2 (BERT-base). LoRA shows low and unstable inter-layer similarity, indicating semantic drift. StructLoRA increases within-block coherence while preserving inter-block diversity.
Appendix K Instruction-tuning evaluation examples
We evaluate instruction-following quality on two representative MT-Bench questions in Fig 7. GPT-4 compares outputs from a baseline LoRA-tuned model and our StructLoRA, consistently rating the latter higher in factual completeness, structure, and engagement.
Instruction-tuning evaluation: StructLoRA enhances reasoning consistency and alignment. Q17: Explain how renewable energy contributes to reducing climate change in a way a middle-school student can understand. LoRA: Renewable energy comes from sources like the sun, wind, and water. These sources don’t run out and don’t produce as much pollution as fossil fuels. Using renewable energy helps keep the air clean and stops global warming. StructLoRA: Renewable energy comes from the sun, wind, and water—things that never run out. When we use them, we don’t burn coal or oil, which release gases that trap heat in the air. Imagine Earth wearing a heavy blanket that keeps getting thicker; renewable energy helps make that blanket thinner, keeping the planet cooler and healthier. GPT-4 Eval: Score 5:8. The baseline is factual but generic. StructLoRA’s answer adds analogy and causal structure, improving clarity and creativity. Q29: Why might some communities resist adopting new technology even when it seems beneficial? LoRA: Some communities may resist new technology because they are used to old ways of doing things. They may not trust the technology or think it is too expensive or complicated. StructLoRA: Communities sometimes hesitate not because they reject progress, but because technology can bring sudden cultural and social changes. For instance, automation may replace jobs or alter traditions. People may fear loss of privacy or local control. Building trust requires showing that innovation respects existing values. GPT-4 Eval: Score 6:8. The baseline is concise but shallow. StructLoRA’s response introduces social, economic, and ethical perspectives, providing balanced reasoning and more context-aware alignment. Summary. StructLoRA produces responses that are more coherent, context-sensitive, and human-aligned. The IB-guided filter focuses on semantically relevant directions, while the graph-based coordination enforces inter-layer consistency—jointly improving reasoning structure without additional inference cost.
Figure 7: Instruction-tuning evaluation on MT-Bench. StructLoRA consistently yields higher GPT-4 evaluation scores than LoRA, generating more structured, informative, and contextually aligned responses across diverse question types.
Appendix L PEFT in Downstream Applications and Future Outlook
Parameter-efficient fine-tuning (PEFT) is now a common way to adapt large pre-trained models to downstream tasks. It is especially useful when full fine-tuning is too expensive or hard to deploy. In recent years, PEFT has moved far beyond language models. It is now widely used in vision, multimodal learning, and domain-specific applications Xin et al. (2024); Zhang et al. (2025c); Xiao et al. (2026b); Liu et al. ; Wang et al. (2025a).
The appeal of PEFT is simple. It lets us reuse strong pre-trained backbones while only updating a small set of task-specific parameters. This makes training cheaper and often more stable. In NLP, this idea has been used for classification, reasoning, and instruction tuning. In vision, PEFT has also shown strong value in real downstream settings, especially when data are limited or domains shift. For example, prompt-based adaptation has been used for cross-domain road damage detection, where the model needs to transfer across different real-world environments Zhang et al. (2025e); Xiao et al. (2026a). Similar ideas have also appeared in scientific and biomedical tasks, where the target domain is specialized and labeled data are often scarce. Prior work on drug–target binding affinity prediction shows that prompt-based adaptation can help connect strong backbone representations with domain-specific downstream objectives Zhang et al. (2025d); Xiao et al. (2024); Zhang et al. (2026). This need for lightweight and transferable adaptation is also common in scientific imaging tasks, where the data distribution is often very different from natural images Xiao et al. (2025a). Besides, recent studies further extend PEFT to data-efficient training paradigms Zhao et al. (2025, 2026). From this view, the main question for PEFT is not only how many parameters to update. A more important question is how to update them well. Our method, StructLoRA, takes a step in this direction. Instead of treating all low-rank directions equally, it asks which directions are useful for the task and how updates across layers should work together. We believe this matters even more in hard downstream settings, such as low-resource learning, cross-domain transfer, and multimodal alignment. In these cases, naive low-rank adaptation may introduce noisy updates or inconsistent behavior across layers.
Looking ahead, PEFT will likely become more structured and more adaptive. One clear direction is to model dependencies across layers instead of tuning each layer in isolation. Another direction is to make the adaptation budget task-aware, so the model can decide where and how much to update based on the downstream objective. PEFT is also likely to play a larger role in scientific and multimodal applications, where data are expensive, models are large, and repeated adaptation is often necessary. As foundation models continue to grow, we expect future PEFT methods to focus not only on parameter efficiency, but also on adaptation quality, robustness, and transferability across tasks and domains.
Appendix M Impact Statement
This work aims to advance the field of machine learning by improving the efficiency and interpretability of parameter-efficient fine-tuning methods. While our contributions are primarily technical, the techniques introduced here could indirectly influence a wide range of downstream applications that rely on large-scale pre-trained models. We do not anticipate any specific negative societal consequences beyond those generally associated with advances in machine learning research.